OpenStudy (anonymous):

40% of what number is 12

OpenStudy (anonymous):

40%=12 40%/3 gives you 10% 10%=3 10% times 10 gives you 100% 100%=30

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

no problem :) you could also do it as: x times 0.4 = 12 x=12/0.4 x=30

OpenStudy (anonymous):

what about 12 % of what number is 105?

OpenStudy (anonymous):

try using the same method :) i'll start you off: 1%=105/12

OpenStudy (anonymous):

Have you done it? :)